# ChronoFlow: A Data-Driven Model for Gyrochronology

### v1

@@ 7 Applications
7.1 New cluster ages
Figure 16: A visualization analogous to Figure 7 showing the probability densities generated by our model at our inferred ages for the Theia 456 stellar stream and the NGC 6709 open cluster, along with the observational data. The stars in NGC 6709 are sized according to cluster membership probability; those values were not provided in the Theia 456 catalog so all stars are sized equally (reflective of our fiducial cluster membership probability of 0.9). We exclude stars having $(BP-RP)_{0}$ 0.3 from our age inference, and have shaded this region in red.
In addition to generating new age estimates for M34, NGC 2516, NGC 1647, and NGC 1750 as described in §5.1, we applied our model to two systems that were not part of our data catalog: the NGC 6709 open cluster (using rotation periods from Cole-Kodikara et al. 2023), and the Theia 456 stellar stream (using rotation periods from Andrews et al. 2022). Gaia EDR3 IDs were provided in both cases, so we calculated the de-reddened photometric uncertainty using the process described in §2.2 using the Edenhofer et al. (2023) dust map. We retained 162 of the 171 Theia 456 stars having reported rotation periods; one was excluded because the ID provided did not match any records in Gaia DR3 or EDR3, and eight more were excluded because they have Gaia RUWE values ¿ 1.4. For NGC 6709, we retained 34 out of the 48 stars with reported rotation periods. 11 were excluded because they were flagged as binaries by Cole-Kodikara et al. (2023), and none of the 37 remaining had a Gaia RUWE ¿ 1.4, but three also had $(BP-RP)_{0}$ values less than 0.3, which we excluded. We then calculated the cluster age posteriors for each, following the methodology described in §5.1 (including the intrinsic scatter in our uncertainties).
@@ 7.1 New cluster ages
Name ChronoFlow Literature Age
Estimate (Myr) Range (Myr)
M34 $241^{+39}_{-33}$ 125–250
NGC 1647 $84^{+24}_{-16}$ 260–365
NGC 1750 $100^{+31}_{-21}$ 250–415
NGC 2516 $119^{+21}_{-18}$ 150–300
NGC 6709 $125^{+43}_{-26}$ 100–190
Theia 456 $142^{+32}_{-20}$ 150–200
@@
We found an age for Theia 456 of $141^{+31}_{-20}$ Myr. Andrews et al. (2022) found ages for Theia 456 (using both isochrones and gyrochronology) of 150-200 Myr, and Kounkel & Covey (2019) estimated it to be $165^{+70}_{-50}$ Myr. Our estimate is on the lower end of these literature values, but consistent at $1\sigma$.
[…]
Figure 16 presents a comparison of the observed data for both systems to the probability density grids from ChronoFlow, analogous to Figure 7. Table 4 summarizes these ages, as well as the four revised cluster ages described in §5.1.1.
Of these six comparisons, the inferred ages of NGC 1647 and NGC 1750 are significantly lower than literature estimates. Both of those clusters have a relatively small sample size in our catalog, and also a dearth of rotators redder than $C_{0}\approx 1.5$, which may contribute to this difference. Nonetheless, Figure 12 shows that our inferred ages do fit ChronoFlow better than the literature ages.
@@ 7.2 Rotational ‘evolution’ tracks
For $(BP-RP)_{0}$ = 0 and 0.5, the expected smooth spindown cause by magnetic braking is not apparent. This may be due to a lack of training data in this regime, and contamination from stars with radiative envelopes (see §5.1).
@@
Examining the reddest stars in our sample, Popinchalk et al. (2021) found that M dwarfs at field ages exhibited a bimodal spin down pattern, with a group of fast rotators having $P_{rot}$ 2 days, and another overdensity at $P_{rot}$ 60 days. Pass et al. (2022) found similar results using wide binaries: some M dwarfs converged onto a slowly rotating sequence quickly, but some remained rotating fast for much longer before spinning down quickly. This bimodal spindown is also a feature of solar mass stars in physical models (e.g. Garraffo et al. 2018). The Garraffo et al. (2018) model has not been tested for fully convective M dwarfs yet, since it relies on convective turnover timescale, which has historically been difficult to measure for such stars. However, Gossage et al. (2024) have recently made progress towards a better method for calculating that parameter. The bimodality with fast and slow rotators is somewhat apparent in our model for $\approx$ solar mass stars, however we do not see the same buildup of fast rotators in the reddest stars. Instead, while we see a narrow distribution of slow rotators appear at $\gtrsim$ 1 Gyr, the distribution of fast rotators is broad.
@@ 7.3 Extrapolation to field ages
Figure 18: Probability densities $\mathcal{P}(P_{rot}\;|\;C_{0},\tau)$ extrapolated to the age of the universe for different characteristic $C_{0}$ values. The vertical line represents are oldest training age (M67). We have plotted the Sun in the third panel for reference, showing that it fits our extrapolated model reasonably well.
@@ 8 Conclusions and Summary
In this work we constructed a rotational catalog of 7,407 stars, from 28 open clusters plus subgroups of the Taurus and Sco-Cen associations. This is the most comprehensive open cluster rotational catalog to date, spanning an age range of 1.5 Myr to 4 Gyr. We have standardized all photometry to Gaia DR3, which we de-reddened using the 3D dustmaps from Edenhofer et al. (2023) and Green et al. (2019). We also propagated uncertainties in parallax, Gaia photometry, and dustmap extinction to get realistic color uncertainties. Finally, we have included cluster membership probabilities, sourced from literature and HDBScan analysis. We present this catalog as part of this work, as well as a deep literature compilation of age estimates for the clusters in the catalog. We hope that this catalog will be a useful benchmark for gyrochronology, and also for the broader astronomical community.
@@
We have shown that ChronoFlow accurately captures the observed $P_{rot}$ dispersion, a result which has eluded many theoretical and empirical models to date. Our model can be used to compute the probability density in color-$P_{rot}$ space at any given age, or examine the $P_{rot}$ density evolution over time for fixed colors. We have also implemented ChronoFlow in a Bayesian framework to infer stellar and cluster age posterior probabilities. The uncertainty on individual stellar ages is $\approx$ 0.7 dex, and we can constrain cluster ages with a statistical uncertainty of $\approx$ 0.06 dex, or 15%. As an example of its applicability, we have used ChronoFlow to calculate new age estimates for five open clusters and one stellar stream.
@@ Appendix B Cuts applied to literature rotation catalogs
      We only include stars with a BinRUWE value of 0. This corresponds to stars with a Gaia RUWE value of ¡1.4, which is a cut we apply in our final selection anyways.
@@
  • We only include stars with a single flag of True and a binary flag of False from Dungee et al. (2022) (Table 2 in that work). We also exclude stars with a false-alarm probablity (FAP) of ¿= 1%, which the authors use as the upper limit for which periodic signals are considered significant. We exclude stars with a measured period of ¿= 175 days, following the authors’ requirement that stars have at least five completed periods within the light-curve duration to be considered reliable.
@@ Appendix C Notes and edge cases from rotation period compilation
C.1 Multiple catalog using the same source measurement
Here is an extended description of the 421 rotation periods that were cited by multiple source catalogs (as described in §2:
@@ C.3 M35
C.3.1 M67
Long et al. (2023) used 10 stars with rotation periods measured by Dungee et al. (2022) that were not considered in the final Dungee et al. (2022) catalog. One of these has a measured rotation period but a FAP ¿ 0.01 (which was used as a quality cut by Dungee et al. 2022). The other nine stars do not actually have rotation periods reported by Dungee et al. (2022) catalog, so we were not able to determine where Long et al. (2023) obtained these values from from. As such, we exclude these 10 stars from our final catalog. The other 45 rotation periods cited by Long et al. (2023) from Dungee et al. (2022) agree with the values presented by the latter.
C.4 NGC2516
@@ Appendix E Cluster membership probabilities
E.2 Douglas et al. (2024): Collinder 135, IC 2391, IC 2602, NGC 2451A, NGC 2547
@@
E.3 Douglas et al. (2019): Hyades and Praesepe
@@
E.4 Rampalli et al. (2021): Praesepe
Rampalli et al. (2021) also compile a membership catalog from multiple sources, and define all as having a ¿70% cluster membership probability. Naively assuming a uniform distribution in probability space, we assign a value of 0.85 for all of these stars.
E.5 NGC 3532
@@ E.10 H Persei
E.12 Curtis et al. (2020): NGC 752, NGC 6819, Pleiades, Ruprecht 147
@@
E.13 Upper Scorpius and $\rho$ Ophiucus subgroups
@@ Appendix F Model Performance Tests
Here we provide further details on the tests we executed to evaluate the performance of ChronoFlow, as described briefly in §5. We used the average stellar log probability as a benchmark metric to measures how well observations align with our $P_{rot}$ density probabilities (e.g., Figures 7 and 8). Averaging the stellar log probabilities accounts for the different sample sizes of each cluster.
@@ G.1 Dustmap comparison
There is a systematic due to the choice of dustmap used to calculate extinctions. Extinction uncertainties are non-trivial due to: (i) parallax (and thus distance) uncertainties, (ii) sampling uncertainties within a single dustmap, and (iii) conversion uncertainties for each photometric band. They are the dominant source of photometric uncertainty in our final data catalog (compared to Gaia systematics and DR3 BP and RP uncertainties). In Figure 24 we present a comparison of $A_{v}$ calculated from the Edenhofer et al. (2023) (E23) and Green et al. (2019) (B19) dustmaps for stars in our catalog that are covered by both. The $A_{V}$ values obtained from the E23 map are systematically lower than those obtained from the B19 map. These results are qualitatively consistent with those found by Diego Godoy-Rivera in a study of Kepler stars covered by both maps (submitted).
@@
The resulting posteriors for each test case are shown in figures 26 and 26. The scatter for all four tests for the Pleiades was small ($\sigma$ = 0.014 dex). The M35 age estimates are more scattered, however the age estimates were consistent when the training dust map was the same as the test dust map. Analyzing the results in a different way, we calculated the scatter in age estimates for both clusters to be 0.01 dex when comparing test cases in which a single dustmap was used. The test cases where dustmaps were mixed resulted in an uncertainty of 0.04 dex. These indicate that the systematic uncertainty due to choice of dustmap is small, if the dustmap is used consistently. However, caution should be taken when applying ChronoFlow to photometry de-reddened with a method other than the E23 or B19 maps.

### v2

@@ 7 Applications
7.1 Ages for coeval populations
Figure 20: A visualization analogous to Figure 7 showing the probability densities generated by our model at our inferred ages for the Theia 456 stellar stream and the NGC 6709 open cluster, along with the observational data. The stars in NGC 6709 are sized according to cluster membership probability; those values were not provided in the Theia 456 catalog so all stars are sized equally (reflective of our fiducial cluster membership probability of 0.9). We exclude stars having $(BP-RP)_{0}<0.5$ from our age inference, and have shaded this region in red.
In addition to generating rotational age estimates for M34, NGC 2516, NGC 1647, and NGC 1750 as described in §5.1, we applied our model to two systems that were not considered in our data catalog: the NGC 6709 open cluster (using rotation periods from Cole-Kodikara et al. 2023), and the Theia 456 stellar stream (using rotation periods from Andrews et al. 2022). Gaia EDR3 IDs were provided in both cases, so we calculated the de-reddened photometric uncertainty using the process described in §2.2 using the Edenhofer et al. (2023) dust map. We retained 161 of the 171 Theia 456 stars having reported rotation periods; one was excluded because the ID provided did not match any records in Gaia DR3 or EDR3, eight more were excluded because they have Gaia RUWE values $>$ 1.4, and one more was excluded because it has $C_{0}<0.5$. For NGC 6709, we retained 34 out of the 48 stars with reported rotation periods. 11 were excluded because they were flagged as binaries by Cole-Kodikara et al. (2023), and none of the 37 remaining had a Gaia RUWE $>$ 1.4, but three also had $C_{0}<0.5$, which we excluded. We then calculated the system age posteriors for each, following the methodology described in §5.1 (including the intrinsic scatter in our uncertainties).
@@ 7.1 Ages for coeval populations
We found an age for Theia 456 of $142^{+26}_{-21}$ Myr. Andrews et al. (2022) found ages for Theia 456 (using both isochrones and gyrochronology) of 150-200 Myr, and Kounkel & Covey (2019) estimated it to be $165^{+70}_{-50}$ Myr. Our estimate is on the lower end of these literature values, but consistent at $1\sigma$, so it supports the previous estimates of Andrews et al. (2022) and Kounkel & Covey (2019).
@@
Figure 20 presents a comparison of the observed data for both systems to the probability density grids from ChronoFlow, analogous to Figure 7.
@@ 7.2 Rotational ‘evolution’ tracks
For $(BP-RP)_{0}$ = 0 and 0.5, the expected smooth spindown caused by magnetic braking is not apparent. This may be due to a lack of training data in this regime, and contamination from stars with radiative envelopes (see §5.1).
@@
Examining the reddest stars in our sample, Popinchalk et al. (2021) found that M dwarfs at field ages exhibited a bimodal spin down pattern, with a group of fast rotators having $P_{rot}<2$ days, and another overdensity at $P_{rot}>60$ days. Pass et al. (2022) found similar results using wide binaries: some M dwarfs converged onto a slowly rotating sequence quickly, but some remained rotating fast for much longer before spinning down quickly. This bimodal spindown is also a feature of solar mass stars in physical models (e.g. Garraffo et al. 2018). The Garraffo et al. (2018) model has not been tested for fully convective M dwarfs yet, since it relies on convective turnover timescale, which has historically been difficult to measure for such stars. However, Gossage et al. (2024) have recently made progress towards a better method for calculating that parameter. The bimodality with fast and slow rotators is somewhat apparent in our model for $\approx$ solar mass stars, however we do not see the same buildup of fast rotators in the reddest stars. Instead, while we see a narrow distribution of slow rotators appear at $\gtrsim$ 1 Gyr, the distribution of fast rotators is broad.
@@ 7.3 Extrapolation to field ages
Figure 22: Probability densities $\mathcal{P}(P_{rot}\;|\;C_{0},\tau)$ extrapolated to the age of the universe for different characteristic $C_{0}$ values. The vertical line represents are oldest training age (M67). We have plotted the Sun in the third panel for reference, showing that it fits our extrapolated model reasonably well. Note that for stars of $\approx$ similar size or larger then the Sun, they will have turned off the MS before the age of the universe, so ChronoFlow would not apply in that regime. This is illustrated qualitatively by the red shaded region, although the exact turnoff age would depend on stellar mass.
@@ 8 Conclusions and Summary
In this work we constructed a rotational catalog of 7,615 stars, from 28 open clusters plus subgroups of the Taurus and Sco-Cen associations. This is the most comprehensive open cluster rotational catalog to date, spanning an age range of 1.5 Myr to 4 Gyr. We have standardized all photometry to Gaia DR3, which we de-reddened using the 3D dustmaps from Edenhofer et al. (2023) and Green et al. (2019). We also propagated uncertainties in parallax, Gaia photometry, and dustmap extinction to get realistic color uncertainties. Finally, we have included cluster membership probabilities, sourced from literature and HDBScan analysis. We present this catalog as part of this work, as well as a deep literature compilation of age estimates for the clusters in the catalog. We hope that this catalog will be a useful benchmark for gyrochronology, and also for the broader astronomical community.
@@
We have shown that ChronoFlow can model the $P_{rot}$ dispersion observed in similar coeval stars across a larger parameter space at open cluster ages than existing literature models (although it does not extend into field ages as GPgyro does). Furthermore, it is at least as accurate as existing models at recovering stellar ages without introducing systematic bias due to model structure. Our model can be used to compute the probability density in color-$P_{rot}$ space at any given age, or examine the $P_{rot}$ density evolution over time for fixed colors. We have also implemented ChronoFlow in a Bayesian framework to infer stellar and cluster age posterior probabilities. The uncertainty on individual stellar ages is $\approx$ 0.7 dex, and we can constrain cluster ages with a statistical uncertainty of $\approx$ 0.06 dex, or 15%. As an example of its applicability, we have used ChronoFlow to calculate new age estimates for five open clusters and one stellar stream.
@@ Appendix B Cuts applied to literature rotation catalogs
      We only include stars with a BinRUWE value of 0. This corresponds to stars with a Gaia RUWE value of $<$1.4, which is a cut we apply in our final selection anyways.
@@
  • We only include stars with a single flag of True and a binary flag of False from Dungee et al. (2022) (Table 2 in that work). We also exclude stars with a false-alarm probablity (FAP) of $\geq$ 1%, which the authors use as the upper limit for which periodic signals are considered significant. We exclude stars with a measured period of $\geq$ 175 days, following the authors’ requirement that stars have at least five completed periods within the light-curve duration to be considered reliable.
@@
  • We exclude stars from the Fritzewski et al. (2020) catalog (NGC 2516) that are included in Table 2 of that work as potential radial velocity binaries.
@@ Appendix C Notes and edge cases from rotation period compilation
C.1 Multiple catalogs using the same source measurement
Here is an extended description of the rotation periods that were cited by multiple source catalogs (as described in §2:
@@ C.1 Multiple catalogs using the same source measurement
  • Both Godoy-Rivera et al. (2021) and Fritzewski et al. (2020) used rotation periods measured by Irwin et al. (2007) for NGC 2516.
@@ C.3 M35
C.4 M67
Long et al. (2023) used 10 stars with rotation periods measured by Dungee et al. (2022) that were not considered in the final Dungee et al. (2022) catalog. One of these has a measured rotation period but a FAP $>$ 0.01 (which was used as a quality cut by Dungee et al. 2022). The other nine stars do not actually have rotation periods reported by Dungee et al. (2022) catalog, so we were not able to determine where Long et al. (2023) obtained these values from from. As such, we exclude these 10 stars from our final catalog. The other 45 rotation periods cited by Long et al. (2023) from Dungee et al. (2022) agree with the values presented by the latter.
C.5 NGC2516
@@ Appendix E Cluster membership probabilities
E.2 Fritzewski et al. (2020): NGC 2516
Fritzewski et al. (2020) considered four membership criteria (photometry, proper motion, parallax, and radial velocity), and only included stars in their rotator catalog that passed at least two of those in their rotator catalog. Since they did not estimate quantitative membership probabilities per star, we therefore assign a default value of 0.9 for all of those stars.
E.3 Douglas et al. (2024): Collinder 135, IC 2391, IC 2602, NGC 2451A, NGC 2547
@@
E.4 Douglas et al. (2019): Hyades and Praesepe
@@
E.5 Rampalli et al. (2021): Praesepe
Rampalli et al. (2021) also compile a membership catalog from multiple sources, and define all as having a $>$70% cluster membership probability. Naively assuming a uniform distribution in probability space, we assign a value of 0.85 for all of these stars.
E.6 NGC 3532
@@ E.11 H Persei
E.13 Curtis et al. (2020): NGC 752, NGC 6819, Pleiades, Ruprecht 147
@@
E.14 Upper Scorpius and $\rho$ Ophiucus subgroups
@@ Appendix F Model Performance Tests
Here we provide further details on the tests we executed to evaluate the performance of ChronoFlow, as described briefly in §5. We used the average stellar log probability as a benchmark metric to measure how well observations aligned with our $P_{rot}$ density probabilities (e.g., Figures 7 and 8). Averaging the stellar log probabilities accounts for the different sample sizes of each cluster.
@@ G.1 Dustmap comparison
There is a systematic due to the choice of dustmap used to calculate extinctions. Extinction uncertainties are non-trivial due to: (i) parallax (and thus distance) uncertainties, (ii) sampling uncertainties within a single dustmap, and (iii) conversion uncertainties for each photometric band. They are the dominant source of photometric uncertainty in our final data catalog (compared to Gaia systematics and DR3 BP and RP uncertainties). In Figure 28 we present a comparison of $A_{v}$ calculated from the Edenhofer et al. (2023) (E23) and Green et al. (2019) (B19) dustmaps for stars in our catalog that are covered by both. The $A_{V}$ values obtained from the E23 map are systematically lower than those obtained from the B19 map. These results are qualitatively consistent with those found by Godoy-Rivera et al. (2025).
@@
The resulting posteriors for each test case are shown in figures 30 and 30. The scatter for all four tests for the Pleiades was small ($\sigma$ = 0.014 dex). The M35 age estimates are more scattered, however the age estimates were consistent when the training dust map was the same as the test dust map. Analyzing the results in a different way, we calculated the scatter in age estimates for both clusters to be 0.01 dex when comparing test cases in which a single dustmap was used. The test cases where dustmaps were mixed resulted in an uncertainty of 0.04 dex. These indicate that the systematic uncertainty due to choice of dustmap is small, if the dustmap is used consistently. However, caution should be taken when applying ChronoFlow to photometry de-reddened with a method other than the E23 or B19 maps.